# Gian Maria Campedelli Fondazione Bruno Kessler

### v1

@@ 1 Introduction
The possibility (and desirability) of reaching human-like AI [^1] Or General Artificial Intelligence or even Superintelligence and AI Singularity, or whatever exotic name associated with AI becoming equally or more intelligent than humans. remains highly debated. And so it has been since 1956, the year in which the Dartmouth Summer Research Project on Artificial Intelligence, the event symbolically marking the beginning of AI as a discipline, was held. Discussions and predictions about human-like AI have been revamped in recent years due to the explosion and diffusion of foundation models, and chiefly Large Language Models (LLMs) (Kim et al.,, 2024; Ishizaki and Sugiyama,, 2025).
@@
Notwithstanding the actual reachability of human-like AI (or the time horizon associated with this scenario), [^2] Admittedly, two topics the author of this piece has no sufficient knowledge to provide definitive answers about. For relevant surveys and reports scanning expert predictions about this very topic, see Müller and Bostrom, (2016); Grace et al., (2018); Association for the Advancement of Artificial Intelligence, (2025). the world – and therefore human society – will soon witness an increasing presence of autonomous AI agents. [^3] While many definitions exist I borrow the popular one proposed by Wooldridge and Jennings, (1995), who wrote that an intelligent or AI agent is a software-based computer system that is characterized by a) autonomy, b) social ability, c) reactivity, and d) pro-activeness. Another broader definition, recently proposed by Mitchell et al., (2025), states that AI agents are “computer software systems capable of creating context-specific plans in non-deterministic environments”. Breakthroughs in intelligent systems have already led to the development and deployment of autonomous agents in different sectors and industries. Prominent examples include the military domain (Palantir,, 2025), finance and banking (Park,, 2024; Bousquette,, 2025), and logistics (Bensinger,, 2025), with the prospect that autonomous AI agents will spread across more and more contexts (e.g., healthcare, see Moritz et al., (2025)).
@@
These developments signal the rise of a hybrid society in which agency is no longer the exclusive prerogative of humans or animals. [^4] Institutions and legal entities also exercise agency. However, they are not central to my argument here, since they can be understood, in a stylized way, as collectives of humans. They are founded and maintained by humans. My focus instead is on entities at the individual level – ontologically, epistemologically, and phenomenologically distinct from humans. Machines fall into this category. AI agents are acquiring capacities to perceive, decide, adapt, and engage socially. This hybridization introduces a novel typology of interactions. For most of history, interaction occurred primarily among biological entities; in recent decades, however, advances in robotics, computing, and especially social media have produced a second modality, centered on human–machine exchanges. This shift has already necessitated new research fields devoted to examining how we communicate, collaborate, and co-exist with technology (Hoc,, 2000; Rahwan et al.,, 2019; Tsvetkova et al.,, 2024).
@@
Nowadays, we stand at the precipice of another paradigm shift, one that may possibly carry consequences of unprecedented scale. The rapid proliferation of truly autonomous (generative) AI agents [^5] I refer to generative AI agents – which are currently the state-of-the-art and may or may not in the future be surpassed by agents built on entirely different premises – as agents powered by foundation models, such as (mostly) LLMs, Vision Foundation Models (VFMs), or Multimodal Models, such as GPT-4o (OpenAI et al.,, 2024). marks the emergence of a third and distinct typology of interaction, i.e., the machine–machine one, a typology that for the first time does not entail any biological entity, one for which our almost complete ignorance may become hugely problematic and consequential (Figure 1).
@@
This critical need for shedding light on machine–machine behavior is already resonating within the AI and computer science communities. Fueled by the widespread use and availability of LLMs, recent scholarship has investigated behavioral patterns of LLM-powered AI agents in different contexts (Dafoe et al.,, 2020; Liu et al.,, 2024; Deng et al.,, 2025; Li et al.,, 2025; Ashery et al.,, 2025). Whether motivated by the potential to simulate complex social phenomena or the desire to understand the emergent dynamics generated by conversations between LLMs, scholars have been attracted by the manifold questions that these new forms of interactions pose for scientific research. In this context, one of the aspects that is fostering notable discussions concerns the risks associated with multi-agent AI systems, i.e., systems of AI agents interacting with each other with no human mediation (Hammond et al.,, 2025; de Witt,, 2025).
@@
Such a discussion is not only speculative and theoretical, but is already substantiated by empirical evidence of unintended deviant and unlawful behaviors by interactive AI agents both in research (Fish et al.,, 2024; Campedelli et al.,, 2024; Bichler et al.,, 2025) as well as in real-world practical domains, as shown by scandals of collusion in algorithmic pricing (Priluck,, 2015). Multi-agent AI systems, in fact, introduce distinct risks by enabling agents to learn from, adapt to, and coordinate with one another in ways that are not always predictable or transparent. This interactive dynamic can give rise to emergent behaviors, patterns of action that are not explicitly programmed and may be difficult to detect, explain, or control. As a result, these systems can generate different types of harm, including fraud, manipulation, discrimination, or the dissemination of disinformation, sometimes absent any direct human intervention and possibly through novel decision-making or atypical behavioral patterns. These developments challenge traditional criminological categories and raise pressing questions about responsibility, regulation, and prevention in a world increasingly shaped by non-human actors.
@@
Inspired by previous theoretical conceptualizations by Woolgar, (1985) and champions of Action-Network Theory (Latour,, 1996; Law and Hassard,, 1999), I contend that we, as a scholarly community, should begin engaging with these foundational issues. I suggest that doing so opens the door to a series of further inquiries, which I will outline and explore in the remainder of this piece. Moreover, I argue that criminologists could contribute – jointly with experts from the AI community – to the efforts to predict, contain, mitigate, and govern the risks emerging from interactive AI agents.
@@ A Tool for Research
Despite the prevailing notion that the convergence of AI and criminology is a recent development, the relationship between these domains traces back to the 1980s (Campedelli,, 2022). Many things have changed since the first attempts to build programs able to predict crime events – not so much in terms of goals, but in terms of popularity, computing power, computational architectures, and richer data availability. In the 1980s and 1990s, attempts at using AI to address crime-related problems were rare and relied on much less powerful hardware, often on symbolic architectures or expert systems (see Icove, (1986), Ratledge and Jacoby, (1989), and Hernandez, (1990)). More recently, the use of machine learning and deep learning has gained traction – with an explosion of publications in the past five years, in criminology and computer science alike. Scholars can now process large amounts of data on personal laptops (or, in the most demanding cases, via cheap cloud servers) and perform prediction or forecasting tasks using more expressive, flexible methods, often based on tree-based or neural architectures.
@@
This compact (and therefore not at all comprehensive) depiction of the current landscape demonstrates that AI in criminology – and, more broadly, in the social sciences – has been seen, studied, and utilized as a tool, a means to an end. In most cases, machine and deep learning algorithms are deployed to solve a specific task (such as forecasting recidivism, e.g., Berk, (2012); Dressel and Farid, (2018), or predictive policing, e.g., Ferguson, (2016), Kaufmann et al., (2019)), to test theories (Molina and Garip,, 2019), or, less commonly, to discover hypotheses in the attempt to generate new research questions in an agnostic fashion, as proposed by Grimmer et al., (2021). In substance, most criminologists and social scientists see AI as a competitor of traditional statistical methods – a set of techniques that make the quantitative researcher’s job easier. [^6] Importantly, the use of AI methods to address criminological research questions applies not only to machine and deep learning approaches but also to LLMs. See, for instance, Adams et al., (2024) and Relins et al., (2025).
@@ A Tool for Committing Crime
AI, unfortunately, is not only seen and utilized as a flexible and powerful tool for research. It is also exploited as a technology for committing crime (Caldwell et al.,, 2020; Blauth et al.,, 2022). King et al., (2020) introduced the term Artificial Intelligence Crime (AIC) to describe the use of AI for unlawful purposes, a phenomenon studied across multiple disciplines. In their seminal paper, they examine three key questions: who should be considered the true perpetrator of an AI-enabled crime (a human or the artificial agent itself?), how an AIC should be defined, and in what ways such crimes are typically carried out.
Building on this, Hayward and Maas, (2021) classify AIC into three subcategories: (1) crimes with AI, (2) crimes against AI, and (3) crimes by AI. The first, and arguably most common, refers to cases where AI is deployed as a tool for malicious purposes, amplifying existing criminal threats and generating new risks. Examples include AI-powered drones used for targeted killings and AI-driven social engineering attacks in cyberspace.
@@
The third category, crimes by AI, encompasses cases where AI operates as an intermediary in unlawful activity. Here, AI’s growing autonomy and capacity for specialized tasks enable it to deviate from deterministic behaviors. Examples include experimental cases of market manipulation and collusion (Martínez-Miranda et al.,, 2016; Ezrachi and Stucke,, 2017), as well as real-world incidents such as an AI agent purchasing illegal goods online (Kasperkevic,, 2015). According to Hayward and Maas, (2021), this subcategory raises critical questions of liability and agency – issues I will return to later in this manuscript.
@@ 2.2.1 AI Agents as an Integral Part of Society
Contemporary AI agents are completely different entities compared to standard Random Forests or Support Vector Machines: the scope of generative AI agents is much broader, characterized by a more diverse set of capabilities and constrained by larger development costs. Notably, all works cited in the previous subsection regarding the use of AI tools for committing crime were published at least four years ago and focused on reinforcement learning approaches rather than generative AI (see Section 3.1 for a discussion of the differences between these two technologies). By contrast, agents powered through generative foundation models emerged recently [^8] A word such as recently has wildly different meanings when comparing the fields of AI and criminology. In the former, the pace of innovation and the sheer volume of publications imply that, in some cases and subfields, work published five years ago is already fatally outdated. In the latter, however, recently may still apply to works published a decade ago or even earlier. I will not elaborate further on this discrepancy, but I am convinced it is related to the broader narrative of this work – namely, the need for criminology and the social sciences to seriously consider how technological breakthroughs may generate societal consequences at a much faster pace than criminology has traditionally accounted for. This point is discussed in detail by Topalli and Nikolovska, (2020). and can communicate, plan, and perceive the environment, solving a multitude of general or specialized tasks with greater speed and versatility than before. [^9] Relevant disclaimer: I am not blind to the many shortcomings of contemporary LLMs, exemplified by (often spectacular) hallucinations and their inability to solve extremely easy problems (Williams and Huckle,, 2024; Xu et al.,, 2025; Malek et al.,, 2025). LLMs (and foundation models in general) have many, clear limits. My argument is not that AI agents are more intelligent than humans; the argument is that they have reached a level of autonomy that allows them to act in interactive environments, that this new collective paradigm requires scholarly attention, and that their failures and hallucinations add a further layer of complexity to understanding and predicting their behaviors. Notably, the argument of this paper is not necessarily tied to generative AI: it would remain relevant even if, in the near future, other technological advancements surpass transformer-based architectures such as LLMs in their cognitive, reasoning, and operational capabilities.
@@ 2.2.2 Theoretical Premises
This call to recognize AI agents as integral social entities is grounded in the fundamental principles of Actor-Network Theory (ANT) (Latour,, 1996; Law and Hassard,, 1999; Latour,, 2007). ANT offers a critical conceptual lens for criminology because it radically flattens the ontological hierarchy between humans and non-humans, which is now essential for understanding the increasing autonomy of AI systems and agents. At its core, ANT conceptualizes all entities – human or non-human, animate or inanimate – as actants of equal analytical importance in the study of society. This perspective deliberately moves away from the assumption that human agency is privileged over the agency of things, including machines. Importantly, and despite the reference to technologies that were very distant from the ones I discuss in this paper, ANT has already been applied in the literature as a social constructivist platform to study and theorize technological advancements and their impact for criminology (Robert and Dufresne,, 2016). Brown, (2006), for instance, argued against a simple binarization separating the human and the artificial, advocating for the use of ANT and the necessity to blend social theory with information theory to really comprehend contemporary criminal phenomena. Aligning with this argument, van der Wagen and Pieters, (2015) studies bot nets, i.e., networks of infected computers controlled by a user, building on the prescriptions of ANT, defining them as hybrid criminal actor-networks, underscoring its relevance to illuminate offending dynamics, victimization as well as countering approaches.
@@
This theoretical perspective requires criminology (and, relatedly, sociology) to re-evaluate the importance of the non-human, echoing the decades-long appeal of Woolgar, (1985). In his seminal work, Woolgar called for a sociology of machines with two specific goals. The first, largely pursued within Science and Technology Studies (STS), concerned the analysis of daily routines and narratives of the AI community and later spurred ethnographic work on algorithmic systems (Seaver,, 2017; Cellard,, 2022; Christin,, 2020), including in criminal justice settings (Brayne and Christin,, 2021). The second goal – less developed but now critical – was precisely to make intelligent machines the actual subject of sociological analysis, challenging the idea that the social is a distinctly human category. As noted by Airoldi, (2021), forty years later Woolgar’s argument is more relevant than ever due to the operational reality of machine agency. Advancements in AI, championed by LLMs and foundation models, make AI agents – technological products equipped with unprecedented computational power and task-solving abilities – available at scale. [^10] Which means, also, that they are available to scholars outside the traditional communities that for decades worked on multi-agent AI systems (see Tan, (1993); Ferber, (1999); Shoham et al., (2007); Sandholm, (2007)), thus enabling broader and more diversified analyses. This technological shift means that what was only possible through abstract theorizing decades ago becomes operationally viable and empirically necessary for criminology today. Scholars can now design and observe the emergent properties of machine-machine networks to anticipate, diagnose or control potential emergent criminal phenomena. Criminologists have not yet ventured into this unexplored path. Yet, scholars in other fields have. Section 3 elaborates on relevant scholarship emerging from the social and computer sciences, with a specific focus on safety. Before that, however, the next subsection provides an operational definition of AI agency, crucial to conceptualize – theoretically and analytically – the entities that are the object of this article.

### v2

@@ 1 Introduction
The possibility (and desirability) of reaching human-like AI [^1] Or General Artificial Intelligence or even Superintelligence and AI Singularity, or whatever exotic name associated with AI becoming equally or more intelligent than humans. remains highly debated. And so it has been since 1956, the year in which the Dartmouth Summer Research Project on Artificial Intelligence, the event symbolically marking the beginning of AI as a discipline, was held. Discussions and predictions about human-like AI have been revamped in recent years due to the explosion and diffusion of foundation models, and chiefly Large Language Models (LLMs) (KimRoadArtificialSuperIntelligence2024; IshizakiLargelanguagemodels2025).
@@
Notwithstanding the actual reachability of human-like AI (or the time horizon associated with this scenario), [^2] Admittedly, two topics the author of this piece has no sufficient knowledge to provide definitive answers about. For relevant surveys and reports scanning expert predictions about this very topic, see MullerFutureProgressArtificial2016c; GraceViewpointWhenWill2018; AssociationfortheAdvancementofArtificialIntelligenceAAAI2025Presidential2025. the world – and therefore human society – will soon witness an increasing presence of autonomous AI agents. [^3] While many definitions exist I borrow the popular one proposed by WooldridgeAgenttheoriesarchitectures1995, who wrote that an intelligent or AI agent is a software-based computer system that is characterized by a) autonomy, b) social ability, c) reactivity, and d) pro-activeness. Another broader definition, recently proposed by MitchellFullyAutonomousAI2025, states that AI agents are “computer software systems capable of creating context-specific plans in non-deterministic environments”. Breakthroughs in intelligent systems have already led to the development and deployment of autonomous agents in different sectors and industries. Prominent examples include the military domain (palantir2025aip), finance and banking (ParkEnhancingAnomalyDetection2024; BousquetteDigitalWorkersHave2025), and logistics (BensingerAmazonsdeliverylogistics2025), with the prospect that autonomous AI agents will spread across more and more contexts (e.g., healthcare, see MoritzCoordinatedAIagents2025).
@@
These developments signal the rise of a hybrid society in which agency is no longer the exclusive prerogative of humans or animals. [^4] Institutions and legal entities also exercise agency. However, they are not central to my argument here, since they can be understood, in a stylized way, as collectives of humans. They are founded and maintained by humans. My focus instead is on entities at the individual level – ontologically, epistemologically, and phenomenologically distinct from humans. Machines fall into this category. AI agents are acquiring capacities to perceive, decide, adapt, and engage socially. This hybridization introduces a novel typology of interactions. For most of history, interaction occurred primarily among biological entities; in recent decades, however, advances in robotics, computing, and especially social media have produced a second modality, centered on human–machine exchanges. This shift has already necessitated new research fields devoted to examining how we communicate, collaborate, and co-exist with technology (Hochumanmachineinteraction2000; RahwanMachinebehaviour2019; Tsvetkovanewsociologyhumans2024b).
@@
Nowadays, we stand at the precipice of another paradigm shift, one that may possibly carry consequences of unprecedented scale. The rapid proliferation of truly autonomous (generative) AI agents [^5] I refer to generative AI agents – which are currently the state-of-the-art and may or may not in the future be surpassed by agents built on entirely different premises – as agents powered by foundation models, such as (mostly) LLMs, Vision Foundation Models (VFMs), or Multimodal Models, such as GPT-4o (OpenAIGPT4oSystemCard2024). marks the emergence of a third and distinct typology of interaction, i.e., the machine–machine one, a typology that for the first time does not entail any biological entity, one for which our almost complete ignorance may become hugely problematic and consequential (Figure 1).
@@
This critical need for shedding light on machine–machine behavior is already resonating within the AI and computer science communities. Fueled by the widespread use and availability of LLMs, recent scholarship has investigated behavioral patterns of LLM-powered AI agents in different contexts (DafoeOpenProblemsCooperative2020; LiuDynamicLLMPoweredAgent2024; DengAIAgentsThreat2025; LiSingleTurnSurveyMultiTurn2025; AsheryEmergentsocialconventions2025). Whether motivated by the potential to simulate complex social phenomena or the desire to understand the emergent dynamics generated by conversations between LLMs, scholars have been attracted by the manifold questions that these new forms of interactions pose for scientific research. In this context, one of the aspects that is fostering notable discussions concerns the risks associated with multi-agent AI systems, i.e., systems of AI agents interacting with each other with no human mediation (HammondMultiAgentRisksAdvanced2025; deWittOpenChallengesMultiAgent2025).
@@
Such a discussion is not only speculative and theoretical, but is already substantiated by empirical evidence of unintended deviant and unlawful behaviors by interactive AI agents both in research (FishAlgorithmicCollusionLarge2024; CampedelliWantBreakFree2024a; BichlerAlgorithmicPricingAlgorithmic2025) as well as in real-world practical domains, as shown by scandals of collusion in algorithmic pricing (PriluckWhenBotsCollude2015). Multi-agent AI systems, in fact, introduce distinct risks by enabling agents to learn from, adapt to, and coordinate with one another in ways that are not always predictable or transparent. This interactive dynamic can give rise to emergent behaviors, patterns of action that are not explicitly programmed and may be difficult to detect, explain, or control. As a result, these systems can generate different types of harm, including fraud, manipulation, discrimination, or the dissemination of disinformation, sometimes absent any direct human intervention and possibly through novel decision-making or atypical behavioral patterns. These developments challenge traditional criminological categories and raise pressing questions about responsibility, regulation, and prevention in a world increasingly shaped by non-human actors.
@@
Inspired by previous theoretical conceptualizations by WoolgarWhynotSociology1985c and champions of Actor-Network Theory (Latouractornetworktheoryfew1996; LawActorNetworkTheory1999), I contend that we, as a scholarly community, should begin engaging with these foundational issues. I suggest that doing so opens the door to a series of further inquiries, which I will outline and explore in the remainder of this piece. Moreover, I argue that criminologists could contribute – jointly with experts from the AI community – to the efforts to predict, contain, mitigate, and govern the risks emerging from interactive AI agents.
@@ A Tool for Research
Despite the prevailing notion that the convergence of AI and criminology is a recent development, the relationship between these domains traces back to the 1980s (CampedelliMachinelearningcriminology2022b). Many things have changed since the first attempts to build programs able to predict crime events – not so much in terms of goals, but in terms of popularity, computing power, computational architectures, and richer data availability. In the 1980s and 1990s, attempts at using AI to address crime-related problems were rare and relied on much less powerful hardware, often on symbolic architectures or expert systems (see IcoveAutomatedCrimeProfiling1986, RatledgeHandbookArtificialIntelligence1989, and HernandezArtificialintelligenceexpert1990). More recently, the use of machine learning and deep learning has gained traction – with an explosion of publications in the past five years, in criminology and computer science alike. Scholars can now process large amounts of data on personal laptops (or, in the most demanding cases, via cheap cloud servers) and perform prediction or forecasting tasks using more expressive, flexible methods, often based on tree-based or neural architectures.
@@
This compact (and therefore not at all comprehensive) depiction of the current landscape demonstrates that AI in criminology – and, more broadly, in the social sciences – has been seen, studied, and utilized as a tool, a means to an end. In most cases, machine and deep learning algorithms are deployed to solve a specific task (such as forecasting recidivism, e.g., BerkCriminalJusticeForecasts2012; Dresselaccuracyfairnesslimits2018d, or predictive policing, e.g., ferguson2016policing, kaufmann2019predictive), to test theories (MolinaMachineLearningSociology2019c), or, less commonly, to discover hypotheses in the attempt to generate new research questions in an agnostic fashion, as proposed by GrimmerMachineLearningSocial2021. In substance, most criminologists and social scientists see AI as a competitor of traditional statistical methods – a set of techniques that make the quantitative researcher’s job easier. [^6] Importantly, the use of AI methods to address criminological research questions applies not only to machine and deep learning approaches but also to LLMs. See, for instance, AdamsNomanshand2024 and RelinsUsingInstructionTunedLarge2025.
@@ A Tool for Committing Crime
AI, unfortunately, is not only seen and utilized as a flexible and powerful tool for research. It is also exploited as a technology for committing crime (CaldwellAIenabledfuturecrime2020; BlauthArtificialIntelligenceCrime2022). KingArtificialIntelligenceCrime2020 introduced the term Artificial Intelligence Crime (AIC) to describe the use of AI for unlawful purposes, a phenomenon studied across multiple disciplines. In their seminal paper, they examine three key questions: who should be considered the true perpetrator of an AI-enabled crime (a human or the artificial agent itself?), how an AIC should be defined, and in what ways such crimes are typically carried out.
Building on this, HaywardArtificialintelligencecrime2021a classify AIC into three subcategories: (1) crimes with AI, (2) crimes against AI, and (3) crimes by AI. The first, and arguably most common, refers to cases where AI is deployed as a tool for malicious purposes, amplifying existing criminal threats and generating new risks. Examples include AI-powered drones used for targeted killings and AI-driven social engineering attacks in cyberspace.
@@
The third category, crimes by AI, encompasses cases where AI operates as an intermediary in unlawful activity. Here, AI’s growing autonomy and capacity for specialized tasks enable it to deviate from deterministic behaviors. Examples include experimental cases of market manipulation and collusion (Martinez-MirandaLearningunfairtrading2016; EzrachiTwoArtificialNeural2017), as well as real-world incidents such as an AI agent purchasing illegal goods online (KasperkevicSwisspolicerelease2015). According to HaywardArtificialintelligencecrime2021a, this subcategory raises critical questions of liability and agency – issues I will return to later in this manuscript.
@@ 2.2.1 AI Agents as an Integral Part of Society
Contemporary AI agents are completely different entities compared to standard Random Forests or Support Vector Machines: the scope of generative AI agents is much broader, characterized by a more diverse set of capabilities and constrained by larger development costs. Notably, all works cited in the previous subsection regarding the use of AI tools for committing crime were published at least four years ago and focused on reinforcement learning approaches rather than generative AI (see Section 3.1 for a discussion of the differences between these two technologies). By contrast, agents powered through generative foundation models emerged recently [^8] A word such as recently has wildly different meanings when comparing the fields of AI and criminology. In the former, the pace of innovation and the sheer volume of publications imply that, in some cases and subfields, work published five years ago is already fatally outdated. In the latter, however, recently may still apply to works published a decade ago or even earlier. I will not elaborate further on this discrepancy, but I am convinced it is related to the broader narrative of this work – namely, the need for criminology and the social sciences to seriously consider how technological breakthroughs may generate societal consequences at a much faster pace than criminology has traditionally accounted for. This point is discussed in detail by TopalliFutureCrimeHow2020. and can communicate, plan, and perceive the environment, solving a multitude of general or specialized tasks with greater speed and versatility than before. [^9] Relevant disclaimer: I am not blind to the many shortcomings of contemporary LLMs, exemplified by (often spectacular) hallucinations and their inability to solve extremely easy problems (WilliamsEasyProblemsThat2024; XuHallucinationInevitableInnate2025; MalekFrontierLLMsStill2025). LLMs (and foundation models in general) have many, clear limits. My argument is not that AI agents are more intelligent than humans; the argument is that they have reached a level of autonomy that allows them to act in interactive environments, that this new collective paradigm requires scholarly attention, and that their failures and hallucinations add a further layer of complexity to understanding and predicting their behaviors. Notably, the argument of this paper is not necessarily tied to generative AI: it would remain relevant even if, in the near future, other technological advancements surpass transformer-based architectures such as LLMs in their cognitive, reasoning, and operational capabilities.
@@ 2.2.2 Theoretical Premises
This call to recognize AI agents as integral social entities is grounded in the fundamental principles of Actor-Network Theory (ANT) (Latouractornetworktheoryfew1996; LawActorNetworkTheory1999; LatourReassemblingSocialIntroduction2007). ANT offers a critical conceptual lens for criminology because it radically flattens the ontological hierarchy between humans and non-humans, which is now essential for understanding the increasing autonomy of AI systems and agents. At its core, ANT conceptualizes all entities – human or non-human, animate or inanimate – as actants of equal analytical importance in the study of society. This perspective deliberately moves away from the assumption that human agency is privileged over the agency of things, including machines. Importantly, and despite the reference to technologies that were very distant from the ones I discuss in this paper, ANT has already been applied in the literature as a social constructivist platform to study and theorize technological advancements and their impact for criminology (RobertActornetworktheorycrime2016). BrowncriminologyhybridsRethinking2006, for instance, argued against a simple binarization separating the human and the artificial, advocating for the use of ANT and the necessity to blend social theory with information theory to really comprehend contemporary criminal phenomena. Aligning with this argument, vanderWagenCybercrimeCyborgCrime2015 studies bot nets, i.e., networks of infected computers controlled by a user, building on the prescriptions of ANT, defining them as hybrid criminal actor-networks, underscoring its relevance to illuminate offending dynamics, victimization as well as countering approaches.
@@
This theoretical perspective requires criminology (and, relatedly, sociology) to re-evaluate the importance of the non-human, echoing the decades-long appeal of WoolgarWhynotSociology1985c. In his seminal work, Woolgar called for a sociology of machines with two specific goals. The first, largely pursued within Science and Technology Studies (STS), concerned the analysis of daily routines and narratives of the AI community and later spurred ethnographic work on algorithmic systems (SeaverAlgorithmsculturetactics2017; CellardAlgorithmsfigurespostdigital2022; Christinethnographeralgorithmblack2020), including in criminal justice settings (BrayneTechnologiesCrimePrediction2021). The second goal – less developed but now critical – was precisely to make intelligent machines the actual subject of sociological analysis, challenging the idea that the social is a distinctly human category. As noted by AiroldiMachineHabitusSociology2021, forty years later Woolgar’s argument is more relevant than ever due to the operational reality of machine agency. Advancements in AI, championed by LLMs and foundation models, make AI agents – technological products equipped with unprecedented computational power and task-solving abilities – available at scale. [^10] Which means, also, that they are available to scholars outside the traditional communities that for decades worked on multi-agent AI systems (see TanMultiagentreinforcementlearning1993; FerberMultiAgentSystemsIntroduction1999; ShohamIfmultiagentlearning2007; SandholmPerspectivesmultiagentlearning2007), thus enabling broader and more diversified analyses. This technological shift means that what was only possible through abstract theorizing decades ago becomes operationally viable and empirically necessary for criminology today. Scholars can now design and observe the emergent properties of machine-machine networks to anticipate, diagnose or control potential emergent criminal phenomena. Criminologists have not yet ventured into this unexplored path. Yet, scholars in other fields have. Section 3 elaborates on relevant scholarship emerging from the social and computer sciences, with a specific focus on safety. Before that, however, the next subsection provides an operational definition of AI agency, crucial to conceptualize – theoretically and analytically – the entities that are the object of this article.